# Eigenstate Typicality as the Dynamical Bridge to the Eigenstate Thermalization Hypothesis: A Derivation from Entropy, Geometry, and Locality

### v1

@@ Abstract
The eigenstate thermalization hypothesis (ETH) provides a powerful framework for understanding thermalization in isolated quantum many-body systems, yet its physical foundations and minimal underlying assumptions remain actively debated. In this work, we develop a unified framework that clarifies the origin of ETH by separating kinematic typicality from dynamical input. We show that the characteristic ETH structure of local operator matrix elements follows from four ingredients: the maximum entropy principle, the geometry of high-dimensional Hilbert space, the locality of physical observables, and a minimal dynamical principle, which we term the eigenstate typicality principle (ETP). ETP asserts that in quantum-chaotic systems, energy eigenstates are statistically indistinguishable from typical states within a narrow microcanonical shell with respect to local measurements. Within this framework, diagonal ETH emerges from measure concentration, while the universal exponential suppression and smooth energy-frequency dependence of off-diagonal matrix elements arise from entropic scaling and local dynamical correlations, without invoking random-matrix assumptions. Our results establish ETH as a consequence of entropy, geometry, and chaos-induced typicality, and clarify its scope, thereby deepening our understanding of quantum thermalization and the emergence of statistical mechanics from unitary many-body dynamics.
@@ I Introduction
where $A_{mn}=\langle E_{m}|A|E_{n}\rangle$, $\omega=E_{m}-E_{n}$, and $E=(E_{m}+E_{n})/2$ are defined in terms of the eigenvalues $E_{m}$ and eigenstates $|m\rangle$ of the many-body Hamiltonian. Here $S(E)$ denotes the thermodynamic entropy, $\mathcal{A}(E)$ is a smooth function giving the microcanonical expectation value of $A$, and $f(E,\omega)$ is a smooth spectral envelope. The factor $R_{mn}$ has zero mean and unit variance and behaves as a pseudo-random variable whose distribution is approximately Gaussian. This structure captures three key ingredients: smooth thermal behavior in the diagonal sector; universal entropic suppression of off-diagonal elements; and a smooth frequency envelope $f(E,\omega)$ governing dynamical correlations, with the remaining fluctuations $R_{mn}$ reflecting quantum-chaotic statistics. ETH has been extensively verified across a broad range of generic nonintegrable quantum many-body systems, including interacting spin chains, fermionic and bosonic lattice models, Floquet systems, and strongly correlated or disorder-averaged random Hamiltonians Rigol2016 ; ETH1 ; ETH2 ; Srednicki1996 ; ETH3 ; ETH4 ; MBL1 ; MBL2 ; MBL3 ; Srednicki99 ; Reimann2016 ; Kim2014 ; Beugeling2013 ; Mondaini2016 ; Jafferis2023 ; Capizzi2025 . These studies have established ETH as a cornerstone of modern quantum statistical mechanics. At the same time, several classes of systems that violate ETH have attracted broad interest, including integrable models integ1 ; integ2 ; integ3 , many-body localized phases MBL1 ; MBL2 ; MBL3 ; Basko2006 ; MBL4 ; MBL5 , many-body critical phases MBC1 ; MBC2 , and quantum many-body scarred systems scar1 ; scar2 ; scar3 ; scar4 ; scar5 . These examples clarify the precise boundaries of ETH and reveal the mechanisms by which thermalization can fail in isolated quantum systems.
@@
Despite extensive evidence, a fully principled derivation of ETH remains incomplete. Existing approaches typically rely on semiclassical arguments rooted in the random-wave model and quantum-classical correspondence ETH2 ; Srednicki1996 ; Srednicki99 ; Berry1977 , on random-matrix assumptions concerning Hamiltonians or eigenvector statistics ETH4 ; HaakeBook ; RMT1 ; RMT2 ; Rigol2016 ; Borgonovi2016 ; Oganesyan2007 ; RMT3 , or on large-scale numerical studies across a wide variety of interacting quantum systems. These methods strongly support ETH but do not identify its minimal microscopic prerequisites. In particular, ETH constrains only the statistics of local observables within a narrow microcanonical shell, whereas assumptions of fully random Hamiltonians or globally random wavefunctions impose far stronger statistical structure than ETH itself requires. Since realistic many-body Hamiltonians are sparse and local, they do not exhibit such excessive randomness. This raises a central question: Can ETH be derived without invoking statistical structures that are stronger than those implied by the ETH ansatz itself?
@@
In this work, we develop a unified entropic framework in which ETH follows from four logically independent ingredients: (i) the maximum entropy principle; (ii) concentration of measure; (iii) locality; and (iv) the eigenstate typicality principle (ETP), which asserts that chaotic eigenstates are typical within a microcanonical shell. The first three are purely kinematic, reflecting the geometry of Hilbert space, while ETP provides the minimal dynamical requirement. Within this framework we obtain both the diagonal and off-diagonal ETH, including the universal $e^{-S(E)/2}$ suppression and the smooth envelopes $\mathcal{A}(E)$ and $f(E,\omega)$, without invoking random matrix assumptions.
@@ II Maximum Entropy, Concentration of Measure, and Canonical Typicality
This section establishes the kinematic framework underlying our entropic approach. We review the maximum-entropy characterization of the microcanonical ensemble, the geometry of narrow energy shells in high-dimensional Hilbert space, and the concentration-of-measure phenomena that underpin canonical typicality Reimann2016 ; Popescu06 ; Goldstein06 ; Lloyd ; Goldstein2010 ; Goldstein20102 ; Reimann2007 ; Linden2009 .
@@
where $\mathbb{E}[f_{A}]$ is the Haar-averaged expectation value of $f_{A}(\psi)$ over the unit sphere in $\mathcal{H}_{E}$, and $\Pr_{\psi}$ denotes probabilities with respect to the same Haar measure. This bound shows that deviations from the microcanonical value are exponentially unlikely in $D$, meaning that almost all pure states in $\mathcal{H}_{E}$ appear thermal to any local observable. Equilibrium behavior thus reflects canonical typicality Reimann2016 ; Popescu06 ; Goldstein06 ; Lloyd ; Goldstein2010 ; Goldstein20102 ; Reimann2007 ; Linden2009 , a geometric property of high-dimensional Hilbert space.
@@
Thus almost all pure states in $\mathcal{H}_{E}$ reproduce the microcanonical expectation value of any local observable. The corresponding variance satisfies (Appendix B)
@@ III Eigenstate Typicality Principle (ETP)
This observation motivates the ETP, which we formulate as follows. For a chaotic many-body Hamiltonian $H$, and any sufficiently narrow microcanonical energy window $\mathcal{H}_{E}$ with dimension $D=\dim(\mathcal{H}_{E})\sim e^{S(E)}\gg 1$, the set of energy eigenstates $\{\,|E_{n}\rangle:E_{n}\in\mathcal{H}_{E}\,\}$ is statistically indistinguishable, for the purposes of local measurements, from states drawn according to the Haar measure on $\mathcal{H}_{E}$, up to corrections that vanish in the thermodynamic limit, for example as $D^{-1}$ or $e^{-S(E)}$.
@@ III.1 Mathematical formulation and hierarchy
$P(\{c_{x}\})\propto\exp\left(-D\sum_{x}|c_{x}|^{2}\right),$
@@ IV.3 Energy and frequency dependence: emergence of $f(E,\omega)$
which separates the universal kinematic suppression dictated solely by Hilbert-space geometry from the system-specific dynamical information carried by $f(E,\omega)$. Under the weak ETP, together with the smooth energy dependence of microcanonical correlation functions, the envelope function $f(E,\omega)$ must itself be smooth. Its dependence on $E$ reflects the thermodynamic continuity of the microcanonical state $\rho_{\mathrm{mc}}(E)$, while its $\omega$-dependence is controlled by the finite-time decay of the connected autocorrelation function $C(t)$ for local operators, so its Fourier transform $S_{A}(\omega)$ cannot have structure on arbitrarily small frequency scales. In chaotic many-body systems, such decay precludes sharp spectral features, ensuring that $f(E,\omega)$ varies smoothly on microscopic energy scales.
@@ IV.4 Summary of ETH derivation
From the derivation of the ETH ansatz, we see that it rests on three foundational pillars that turn thermalization into a precise statement about energy eigenstates: 1. Canonical Typicality. Generic pure states within a narrow microcanonical shell $\mathcal{H}_{E}$ reproduce microcanonical expectation values for all $k$-local observables, with fluctuations exponentially small in the shell dimension $D$. This is a purely geometric property of high-dimensional Hilbert spaces, originating from concentration of measure and independent of any Hamiltonian dynamics. 2. ETP. ETP provides the essential dynamical input. In quantum-chaotic systems, energy eigenstates themselves belong to this typical set: for local observables, they are statistically indistinguishable from Haar-random states drawn from $\mathcal{H}_{E}$, at least at the level of low-order moments (weak ETP). In this way, kinematic typicality results are transferred into the Hamiltonian eigenbasis. 3. Locality and sparse operator support. A $k$-local operator $A$ probes only a vanishingly small sector of the exponentially large Hilbert space. This locality fixes the universal ETH suppression factor $e^{-S(E)/2}=D^{-1/2}$: the geometry of the microcanonical shell yields the $D^{-1/2}$ scaling of typical fluctuations, while locality ensures $\mathrm{Tr}(\Pi_{E}A^{2})=O(D)$. Consequently, the entropy suppression in the off-diagonal ETH is ultimately a consequence of the entropic geometry of Hilbert space and the local measurements. Locality further implies the smoothness of the ETH envelope functions. The diagonal function $\mathcal{A}(E)$ inherits its smoothness from the thermodynamic continuity of the microcanonical ensemble, while the envelope $f(E,\omega)$ is smooth because local dynamics exhibit finite-time decay of correlations. This behavior is consistent with Lieb-Robinson bounds, which preclude arbitrarily sharp spectral features in $S_{A}(\omega)$ (see Appendix E).
@@ V Conclusion and Discussion
Dynamics enters only through the Eigenstate Typicality Principle (ETP), which states that chaotic many-body eigenstates populate the energy shell as typical vectors. ETP is far weaker than assuming a random-matrix Hamiltonian; it constrains only eigenvectors, not the distribution of matrix elements in the Hamiltonian itself. Once ETP holds, the ETH structure follows naturally. The diagonal elements equal the microcanonical average because typical states have sharply concentrated local properties, and off-diagonal elements are exponentially suppressed because local operators occupy only a vanishing fraction of the global Hilbert space. Locality then enforces the smooth frequency dependence of $f(E,\omega)$. Thus, thermalization emerges from the interplay between kinematics of high-dimensional state space and the ability of chaotic dynamics to explore that space. Chaos does not generate thermality; it renders the system typical enough for kinematic constraints to dictate thermodynamic behavior.
Future directions follow naturally from this entropic perspective. A central challenge is to understand how eigenstates evolve from atypical to typical as integrability is gradually broken, and how this crossover is reflected in entanglement scaling, operator complexity, and the structure of eigenstate spectra. Another promising direction is to extend this framework to constrained Hilbert spaces, where additional physical constraints restrict the accessible microcanonical manifold. In such systems, altered state counting and modified Hilbert space geometry may reshape the entropy $S(E)$ and lead to distinct forms of ETH, offering new insight into thermalization beyond conventional lattice models. For Floquet systems, where energy is not conserved, the relevant “energy shell” is the unitarily evolving Hilbert space of the Floquet operator, and the ETP therefore predicts a natural infinite-temperature form of ETH. In open systems governed by Lindblad dynamics, it remains an open question whether a generalized typicality applies to non-equilibrium steady states and how dissipation modifies the characteristic entropic suppression of matrix elements. Connections to quantum information provide a further avenue, where entropic constraints may be probed through circuit complexity, scrambling diagnostics, or entanglement spectroscopy. These tools suggest that chaotic typicality can be characterized not only through thermodynamic behavior but also through the scaling of informational resources, pointing toward a broader unification of thermalization, quantum information, and complexity.
@@ Acknowledgements.
For a $k$-local operator $A$, we define its matrix elements within the energy shell as $A_{ij}=\langle i|A|j\rangle$.
In Eq. (8) of the main text, we evaluated the Haar expectation of $f_{A}(\psi)$ We now present the detailed derivation of its variance, $\mathrm{Var}(f_{A})=\mathbb{E}[f_{A}(\psi)^{2}]-\left(\mathbb{E}[f_{A}(\psi)]\right)^{2}$.
@@
$S_{A}(\omega)\propto\sum_{m\neq n}\delta(\omega-\omega_{mn})|A_{mn}|^{2}.$
Appendix Appendix E Localization of spectral weight from Lieb-Robinson bounds
The smoothness of the envelope $f(E,\omega)$ in the frequency domain is not an additional assumption but a direct consequence of locality, as embodied by the Lieb-Robinson bound.
For local operators $A$ and $B$ in a gapped or gapless system with short-range interactions, the Lieb-Robinson bound states:
@@
Now, consider the autocorrelation function $C(t)=\langle A(t)A(0)\rangle_{\mathrm{mc}}^{\mathrm{conn}}$. The operator $A(t)$ spreads spatially, but its support remains within a light-cone of size $\sim v_{\mathrm{LR}}t$. For a local operator $A(0)$ at the origin, $C(t)$ will decay (e.g., algebraically or exponentially) for times $|t|\gtrsim t_{*}$, where $t_{*}$ is a characteristic relaxation time that is independent of system size $L$. This is a fundamental property of local dynamics in themodynamic systems.
The spectral function $S_{A}(\omega)$ is the Fourier transform of $C(t)$:
$S_{A}(\omega)=\int_{-\infty}^{\infty}dt\,e^{-i\omega t}C(t).$
The finite relaxation time $t_{*}$ implies that $S_{A}(\omega)$ is a smooth function of $\omega$ on scales larger than $\Delta\omega\sim 1/t_{*}$. It cannot have sharp features or fine structure on arbitrarily small frequency scales. Since $|f(E,\omega)|^{2}$ is directly proportional to $S_{A}(\omega)$ (up to the density of states factor), it must inherit this smoothness.

### v2

# Deriving the Eigenstate Thermalization Hypothesis from Eigenstate Typicality and Kinematic Principles

@@ Abstract
The eigenstate thermalization hypothesis (ETH) provides a powerful framework for understanding thermalization in isolated quantum many-body systems, yet a complete and conceptually transparent derivation has remained elusive. In this work, we derive the structure of ETH from a minimal dynamical principle, which we term the eigenstate typicality principle (ETP), together with general kinematic ingredients arising from entropy maximization, Hilbert-space geometry, and locality. ETP asserts that in quantum-chaotic systems, energy eigenstates are statistically indistinguishable, with respect to local measurements, from states drawn from the Haar measure on a narrow microcanonical shell. Within this framework, diagonal ETH arises from concentration of measure, provided that eigenstate typicality holds. The structure of off-diagonal matrix elements is then fixed by entropic scaling and the finite-time dynamical correlations of local observables, with ETP serving as the dynamical bridge to energy eigenstates, without invoking random-matrix assumptions. Our results establish ETH as a consequence of entropy, Hilbert-space geometry, and chaos-induced eigenstate typicality, and clarify its regime of validity across generic quantum-chaotic many-body systems, thereby deepening our understanding of quantum thermalization and the emergence of statistical mechanics from unitary many-body dynamics.
@@ I Introduction
where $A_{mn}=\langle E_{m}|A|E_{n}\rangle$, $\omega=E_{m}-E_{n}$, and $E=(E_{m}+E_{n})/2$ are defined in terms of the eigenvalues $E_{m}$ and eigenstates $|m\rangle$ of the many-body Hamiltonian. Here $S(E)$ denotes the thermodynamic entropy, $\mathcal{A}(E)$ is a smooth function giving the microcanonical expectation value of $A$, and $f(E,\omega)$ is a smooth spectral envelope. The factor $R_{mn}$ has zero mean and unit variance and behaves as a pseudo-random variable whose distribution is approximately Gaussian. This structure captures three key ingredients: smooth thermal behavior in the diagonal sector; universal entropic suppression of off-diagonal elements; and a smooth frequency envelope $f(E,\omega)$ governing dynamical correlations, with the remaining fluctuations $R_{mn}$ reflecting quantum-chaotic statistics. ETH has been extensively verified across a broad range of generic nonintegrable quantum many-body systems, including interacting spin chains, fermionic and bosonic lattice models, Floquet systems, and strongly correlated or disorder Hamiltonians Rigol2016 ; ETH1 ; ETH2 ; Srednicki1996 ; ETH3 ; ETH4 ; MBL1 ; MBL2 ; MBL3 ; Srednicki99 ; Reimann2016 ; Kim2014 ; Beugeling2013 ; Mondaini2016 ; Jafferis2023 ; Capizzi2025 . These studies have established ETH as a cornerstone of modern quantum statistical mechanics. At the same time, several classes of systems that violate ETH have attracted broad interest, including integrable models integ1 ; integ2 ; integ3 , many-body localized phases MBL1 ; MBL2 ; MBL3 ; Basko2006 ; MBL4 ; MBL5 , many-body critical phases MBC1 ; MBC2 , and quantum many-body scarred systems scar1 ; scar2 ; scar3 ; scar4 ; scar5 . These examples clarify the precise boundaries of ETH and reveal the mechanisms by which thermalization can fail in isolated quantum systems.
@@
Despite extensive evidence, a fully transparent and minimal derivation of ETH remains incomplete. Existing approaches typically rely on semiclassical arguments rooted in the random-wave model and quantum-classical correspondence ETH2 ; Srednicki1996 ; Srednicki99 ; Berry1977 , on random-matrix assumptions concerning Hamiltonians or eigenvector statistics ETH4 ; HaakeBook ; RMT1 ; RMT2 ; Rigol2016 ; Borgonovi2016 ; Oganesyan2007 ; RMT3 , or on large-scale numerical studies across a wide variety of interacting quantum systems. These methods strongly support ETH but do not identify its minimal microscopic prerequisites. In particular, ETH constrains only the statistics of local observables within a narrow microcanonical shell, whereas assumptions of fully random Hamiltonians or globally random wavefunctions impose far stronger statistical structure than ETH itself requires. Since realistic many-body Hamiltonians are sparse and local, they do not exhibit such excessive randomness. This raises a central question: Can ETH be derived without invoking statistical structures that are stronger than those implied by the ETH ansatz itself?
@@
In this work, we develop a unified entropic framework in which ETH follows from two conceptually distinct ingredients. The first is a set of general kinematic principles, including the maximum entropy characterization of the microcanonical ensemble, concentration of measure in high-dimensional Hilbert spaces, and the locality of few-body operators. These ingredients reflect the geometric and entropic structure of Hilbert space and are entirely independent of dynamics. The second is a minimal dynamical principle, which we term the eigenstate typicality principle (ETP), asserting that chaotic energy eigenstates are typical within a narrow microcanonical shell with respect to local measurements. Within this framework, we obtain both diagonal and off-diagonal ETH, including the universal $e^{-S(E)/2}$ suppression and the smooth envelopes $\mathcal{A}(E)$ and $f(E,\omega)$, without invoking random-matrix assumptions.
@@ II Maximum Entropy, Concentration of Measure, and Canonical Typicality
This section establishes the kinematic framework underlying our entropic approach. We review the maximum-entropy characterization of the microcanonical ensemble, the geometry of narrow energy shells in high-dimensional Hilbert space, and the concentration-of-measure phenomena that underpin canonical typicality Reimann2016 ; Popescu06 ; Goldstein06 ; Lloyd ; Goldstein2010 ; Goldstein20102 ; Reimann2007 ; Linden2009 ; Yang2015 .
@@
where $C>0$ is a universal constant, $\mathbb{E}[f_{A}]$ is the Haar-averaged expectation value of $f_{A}(\psi)$ over the unit sphere in $\mathcal{H}_{E}$, and $\Pr_{\psi}$ denotes probabilities with respect to the same Haar measure. We will demonstrate below that $\mathbb{E}[f_{A}]$ coincides with the microcanonical expectation value of the operator $A$. Thus, this bound shows that deviations from the microcanonical value are exponentially unlikely in $D$, meaning that almost all pure states in $\mathcal{H}_{E}$ appear thermal to any local observable. Equilibrium behavior thus reflects canonical typicality Reimann2016 ; Popescu06 ; Goldstein06 ; Lloyd ; Goldstein2010 ; Goldstein20102 ; Reimann2007 ; Linden2009 , a geometric property of high-dimensional Hilbert space.
@@
Here the subscript in $\mathbb{E}_{\psi}[\cdot]$ emphasizes that the average is taken over random pure states $\psi$ within the microcanonical shell, which is the same ensemble underlying the expectation value in Eq. (7). Thus almost all pure states in $\mathcal{H}_{E}$ reproduce the microcanonical expectation value of any local observable. The corresponding variance satisfies (Appendix B)
@@ III Eigenstate Typicality Principle (ETP)
This observation motivates the ETP, which we formulate as follows. For a chaotic many-body Hamiltonian $H$, and any sufficiently narrow microcanonical energy window $\mathcal{H}_{E}$ with dimension $D\gg 1$, the set of energy eigenstates $\{\,|E_{n}\rangle:E_{n}\in\mathcal{H}_{E}\,\}$ is statistically indistinguishable, for the purposes of local measurements, from states drawn according to the Haar measure on $\mathcal{H}_{E}$, up to corrections that vanish in the thermodynamic limit, for example as $D^{-1}$ or $e^{-S(E)}$.
@@ III.1 Mathematical formulation and hierarchy
$P(\{c_{x}\})\propto\exp\left(-D\sum_{x}|c_{x}|^{2}\right),$ (15)
@@ IV.3 Energy and frequency dependence: emergence of $f(E,\omega)$
which separates the universal kinematic suppression dictated solely by Hilbert-space geometry from the system-specific dynamical information carried by $f(E,\omega)$. Under the weak ETP, together with the smooth energy dependence of microcanonical correlation functions, the envelope function $f(E,\omega)$ must itself be smooth. Its dependence on $E$ reflects the thermodynamic continuity of the microcanonical state $\rho_{\mathrm{mc}}(E)$, while its $\omega$-dependence is controlled by the finite-time decay of the connected autocorrelation function $C(t)$ for local operators (see Appendix D), so its Fourier transform $S_{A}(\omega)$ cannot have structure on arbitrarily small frequency scales. In chaotic many-body systems, such decay precludes sharp spectral features, ensuring that $f(E,\omega)$ varies smoothly on microscopic energy scales.
@@ IV.4 Summary of ETH derivation
From the derivation of the ETH ansatz, we see that it rests on three foundational pillars that turn thermalization into a precise statement about energy eigenstates: 1. Canonical Typicality. Generic pure states within a narrow microcanonical shell $\mathcal{H}_{E}$ reproduce microcanonical expectation values for all $k$-local observables, with fluctuations exponentially small in the shell dimension $D$. This is a purely geometric property of high-dimensional Hilbert spaces, originating from concentration of measure and independent of any Hamiltonian dynamics. 2. ETP. ETP provides the essential dynamical input. In quantum-chaotic systems, energy eigenstates themselves belong to this typical set: for local observables, they are statistically indistinguishable from Haar-random states drawn from $\mathcal{H}_{E}$, at least at the level of low-order moments (weak ETP). In this way, kinematic typicality results are transferred into the Hamiltonian eigenbasis. ETP reflects the ergodic and unbiased nature of the distribution of a system’s energy eigenstates. 3. Locality and sparse operator support. A $k$-local operator $A$ probes only a vanishingly small sector of the exponentially large Hilbert space. This locality fixes the universal ETH suppression factor $e^{-S(E)/2}=D^{-1/2}$: the geometry of the microcanonical shell yields the $D^{-1/2}$ scaling of typical fluctuations, while locality ensures $\mathrm{Tr}(\Pi_{E}A^{2})=O(D)$. Consequently, the entropy suppression in the off-diagonal ETH is ultimately a consequence of the entropic geometry of Hilbert space and the local measurements. Locality further implies the smoothness of the ETH envelope functions. The diagonal function $\mathcal{A}(E)$ inherits its smoothness from the thermodynamic continuity of the microcanonical ensemble, while the envelope $f(E,\omega)$ is smooth because local dynamics exhibit finite-time decay of correlations.
@@ V Conclusion and Discussion
Dynamics enters only through the Eigenstate Typicality Principle (ETP), which states that chaotic many-body eigenstates populate the energy shell as typical vectors. ETP is far weaker than assuming a random-matrix Hamiltonian; it constrains only eigenvectors, not the distribution of matrix elements in the Hamiltonian itself. Once ETP holds, the ETH structure follows naturally. The diagonal elements equal the microcanonical average because typical states have sharply concentrated local properties, and off-diagonal elements are exponentially suppressed because local operators occupy only a vanishing fraction of the global Hilbert space. Locality then enforces the smooth frequency dependence of $f(E,\omega)$. Thus, thermalization emerges from the interplay between kinematics of high-dimensional state space and the ability of chaotic dynamics to explore that space. The role of quantum chaos is not to determine the specific structure of thermal behavior, but to ensure eigenstate typicality, thereby allowing entropic and kinematic constraints to govern thermalization.
Future directions follow naturally from this entropic perspective. A central challenge is to understand how eigenstates evolve from atypical to typical as integrability is gradually broken, and how this crossover is reflected in entanglement scaling, operator complexity, and the structure of eigenstate spectra. Another promising direction is to extend this framework to constrained Hilbert spaces, where additional physical constraints restrict the accessible microcanonical manifold. In such systems, altered state counting and modified Hilbert space geometry may reshape the entropy $S(E)$ and lead to distinct forms of ETH, offering new insight into thermalization beyond conventional lattice models. In open systems governed by Lindblad dynamics, it remains an open question whether a generalized typicality applies to non-equilibrium steady states and how dissipation modifies the characteristic entropic suppression of matrix elements. Connections to quantum information provide a further avenue, where entropic constraints may be probed through circuit complexity, scrambling diagnostics, or entanglement spectroscopy. These tools suggest that chaotic typicality can be characterized not only through thermodynamic behavior but also through the scaling of informational resources, pointing toward a broader unification of thermalization, quantum information, and complexity.
@@ Acknowledgements.
For a $k$-local operator $A$, we define its matrix elements within the energy shell as $A_{ij}=\langle i|A|j\rangle$. In Eq. (8) of the main text, we evaluated the Haar expectation of $f_{A}(\psi)$ We now present the detailed derivation of its variance, $\mathrm{Var}(f_{A})=\mathbb{E}[f_{A}(\psi)^{2}]-\left(\mathbb{E}[f_{A}(\psi)]\right)^{2}$.
@@
$S_{A}(\omega)\propto\frac{1}{D}\sum_{m\neq n}\delta(\omega-\omega_{mn})|A_{mn}|^{2}.$
The smoothness of the envelope $f(E,\omega)$ in the frequency domain is not an additional assumption but a direct consequence of locality, as embodied by the Lieb-Robinson bound. For local operators $A$ and $B$ in a gapped or gapless system with short-range interactions, the Lieb-Robinson bound states:
@@
Now, consider the autocorrelation function $C(t)$. The operator $A(t)$ spreads spatially, but its support remains within a light-cone of size $\sim v_{\mathrm{LR}}t$. For a local operator $A(0)$ at the origin, $C(t)$ will decay (e.g., algebraically or exponentially) for times $|t|\gtrsim t_{*}$, where $t_{*}$ is a characteristic relaxation time that is independent of system size $L$. This is a fundamental property of local dynamics in themodynamic systems. The finite relaxation time $t_{*}$ implies that the spectral function $S_{A}(\omega)$ is a smooth function of $\omega$ on scales larger than $\Delta\omega\sim 1/t_{*}$. It cannot have sharp features or fine structure on arbitrarily small frequency scales. Since $|f(E,\omega)|^{2}$ is directly proportional to $S_{A}(\omega)$ (up to the density of states factor), it must inherit this smoothness.